Consider the following reaction at 298K.
I2 (s) + 2 Fe2+ (aq) 2 I- (aq) + 2 Fe3+ (aq)

Which of the following statements are correct?

Choose all that apply.
ΔGo > 0
K < 1
The reaction is product-favored.
n = 2 mol electrons
Eocell > 0


Submit Answer
1
Expert's answer
2019-04-17T04:15:29-0400
"E_{cell}^0 = E_{Red}^0 - E_{Ox}^0"

"I_2(s) + 2e^- \rightarrow 2I^-" "E_{Red}^0 = 0.54 V"

"(Fe^{2+} \rightarrow Fe^{3+} +1e^-)\times2" "E_{Ox}^0=-0.77 V"

"E_{cell}^0= 0.54+(-0.77)= -0.23 V"

"\Delta G_0= -nFE_{cell}^0= -2\times 96485.309\times (-0.23)= 44383\frac{J}{mol}"

"\Delta G = -RTlnK"

"44383= -8.314\times 8.314lnK"

"K= e^{-\frac{44383}{8.314\times298}}= 1.66\times10^{-8}"

Right answers are:

"\Delta G_0>0"

"K<1"